OpenStudy (anonymous):

The slope of a line is –2 and its y-intercept is (0, 3). What is the equation of the line that is parallel to the first line and passes through (2, 2)?

OpenStudy (anonymous):

A. 2x + y = 6 B. y = –2x + 3 C.y=1/2x+6 D. y = –2x – 6